OpenStudy (anonymous):

if a,b,c are three given coplanar vectors such that a=kb+jc .k and j are scalars then find k and j?

OpenStudy (anonymous):

its hard!

OpenStudy (perl):

so `a` is a linear combination of vectors `b` and `c`

OpenStudy (loser66):

 if they are as shown, then k = j =1

OpenStudy (anonymous):

If they are coplanar, then you an treat them like 2D vectors

OpenStudy (anonymous):

But you only have two equations and like, eight variables, so you need more info.

ganeshie8 (ganeshie8):

essentially the question is asking you to solvea full column rank matrix when n=2 : \[\begin{bmatrix} b_1& c_1\\b_2&c_2\\b_3&c_3\\\cdots\\\end{bmatrix}\begin{bmatrix}k\\j\end{bmatrix} = \begin{bmatrix} a_1\\a_2\\a_3\\\cdots \end{bmatrix}\]

ganeshie8 (ganeshie8):

assuming ofcourse b and c are heading in different directions then you will have an unique solution in 2D this solution equals [b c]^(-1) a

ganeshie8 (ganeshie8):

existence of unique solution is easy to see for a full column matrix as the right hand side is in columnspace of the left side matrix

OpenStudy (akonkel):

@danish071996